### v1

# Capital Games and Growth Equilibria

Abstract
We examine formal games that we call ”capital games” in which player payoffs are known, but their payoffs are not guaranteed to be von Neumann-Morgenstern utilities. In capital games, the dynamics of player payoffs determine their utility functions. Different players can have different payoff dynamics. We make no assumptions about where these dynamics come from, but implicitly assume that they come from the players’ actions and interactions over time. We define an equilibrium concept called ”growth equilibrium” and show a correspondence between the growth equilibria of capital games and the Nash equilibria of standard games.
1 Background on Utility Functions
Why is the use of expectation values and utility functions so ubiquitous in game theory, mechanism design, auction theory, decision theory, and economics more broadly? Part of the reason may be that in their seminal book “The Theory of Games and Economic Behavior”  Von Neumann and Morgenstern (VNM) proved what is called the Von Neumann–Morgenstern Utility Theorem.
Suppose an individual is presented with a set of lotteries, where each lottery is a probability distribution over a set of possible outcomes The VNM Utility Theorem is a statement about preferences over lotteries, where a lottery is a probability distribution over a set of possible outcomes. The theorem tells us that for any preference over lotteries that satisfies basic axioms, there is a way to assign real values, or “utilities”, to the outcomes of the lotteries such that the lotteries’ expectation values characterize the preference relation.
VNM Utility Theorem
Define a lottery to be a probability distribution over a set. We will denote that lottery $A$ is strictly preferred to lottery $B$ by $A\succ B$, indifference between the lotteries as $A\sim B$, and weak preference by $A\succeq B$. We now define four axioms that characterize preference relations.
  • Completeness: For every pair of lotteries, either $A\succeq B$ or $B\succeq A$
  • Transitivity: If $A\succeq B$ and $B\succeq C$ then $A\succeq C$
  • Continuity: If $A\succeq B\succeq C$ then there exists a probability $p\in[0,1]$ such that $pA+(1-p)C\sim B$
  • Independence: For any lottery $C$ and probability $p$, $A\succeq B$ if and only if $(1-p)A+pC\succeq(1-p)B+pC$
Theorem 1 (VNM Utility Theorem).
For any preference relation satisfying Completeness, Transitivity, Continuity, and Independence, there is a utility function $u$ that assigns a real number to every possible outcome of the lotteries such that for any two lotteries $A,B$, $A\succ B$ if and only if $E[u(A)]>E[u(B)]$.
Since “utility” is completely abstract, and the units do not have inherent physical meaning, this is an entirely general result. Similarly, lotteries are defined over any set, so the elements of the set that serves as outcomes do not need to have any natural real value associated with them. VNM left open the questions of (1) what our preference relations over lotteries should be in any given setting, and (2) how to determine what VNM utility function to use in any given situation.
Confusing Observables with Utilities
An observable is a mapping of lottery outcomes to real values. For any preference relation over lotteries there are an infinite number of observables that are not VNM utilities, because their expectation values to not characterize the preference relation. VNM utility is defined such that we want to maximize its expected value, but that is not the case for all observables. To quote VNM, “We have practically defined numerical utility as being that thing for which the calculus of mathematical expectations is legitimate.”-Von Neumann and Morgenstern (2007). It is a remarkably common mistake to take an observable and treat it as though it is necessarily a VNM utility function, inferring a preference relation from its expectation values. We will call this mistake the Fallacy of Utility Conflation.
Definition 1 (Fallacy of Utility Conflation).
The fallacy of utility conflation occurs when assuming that an observable must be a VNM utility function, and that its expectation values necessarily induce the appropriate ordering over lotteries.
Preferences Over Observables Do Not Imply Utilities
The Fallacy of Utility Conflation (FUC Fallacy) appears most commonly where the outcomes of a lottery are already real-valued observables, e.g., monetary payments. When the values of the observable are something an individual would like to be greater, i.e. a “good”, it is easy to mistakenly believe that the observable is a VNM utility. This confuses the use of “utility” in philosophical utilitarianism with a VNM utility. If one can choose an outcome deterministically you may want the outcome with the largest value, but for two different lotteries over those outcomes you do not necessarily want the lottery that maximizes the expected value of the outcome.
For a given a set of lotteries with real-valued outcomes, there is no singular preference relation over the lotteries or VNM utility function implied by them, even if there is a clear preference relation over the outcomes themselves. We are still faced with the same questions of what our preference relations over lotteries should be and how to determine what VNM utility function to use in any given situation. Lotteries over observables, with or without preferences over the lottery outcomes, are not enough on their own to give us an answer. The problem of what our preference should be over lotteries is still under-determined, and this is frequently misunderstood.
Example: The Infamous Coin Flip
Peters and Gell-Mann (2016) provide a prime example of a lottery with real-valued outcomes, for which an individual has a clear preference over the outcomes, and yet this does not determine a VNM utility function. The lottery outcome is determined by the flip of a fair coin. The observable is money – you get a certain amount of money based on the outcome of the coin flip. While you are assumed to always want more money, that doesn’t mean that your VNM utility is the monetary payoff.
Deducing VNM Utility Functions from Dynamics
One way to derive VNM utility functions from lotteries over real-values is to assume that the values experiences dynamics – changes over time. In other words, you treat each decision of choosing between lotteries as if the lottery chosen will be one step in a larger stochastic process. The assumptions made about the stochastic process, or uncertainty about the process, underlie how the choice between lotteries determines the dynamics of the observable. Rather than choosing between lotteries, we have lifted the problem to become preferences over stochastic processes which are realized over time. To choose between stochastic processes we need a single additional assumption; a choice axiom.
Growth Rates and Ergodicity
Ergodicity Economics provides an axiom for choosing between stochastic processes; choose the stochastic process with the greatest time-average growth rate. Growth rates are ergodic, meaning that we can maximize their time-average by maximizing their expected value. Hence, for any stochastic process capturing the dynamics of a desirable observable, it’s growth rate gives us the analog of a VNM utility function, except that it characterizes preferences over stochastic processes rather than preferences over lotteries. We now have a fairly general tool for constructing VNM utility functions when outcomes are real-valued and realized over time. When faced with a choice between real-valued lotteries that are realized over time, we can imagine the lotteries each as one step in a stochastic process, and derive our VNM utility function from their respective growth rates. Naturally, this requires us to make assumptions about the stochastic processes of which each lottery is a part.
Games and Best Responses
In non-cooperative game theory, every player’s “best response” to the strategies of other players is the strategy that maximizes the expected value of their utility. However, maximizing expected value is only the right decision criteria for choosing among strategies when their payoff is expressed in units of VNM utility. Given any real-valued observable for game outcomes, which may not be a VNM utility, we can use growth rates to find the VNM utility function whose expectation value we want to maximize as our “best response.”
Contributions
We introduce a different model than is typically used in standard game theory and mechanism design. We create what we call capital games, where payoffs are an observable in units of “capital” which are not assumed to be the VNM utilities of each player. Given certain assumptions about capital dynamics, we can determine players’ “best responses.” We endow each player in the game with an initial amount of capital in all capital games. For capital games, we show that equilibria always exist when players try to maximize the time-average growth rate of their wealth, as a generalization of Nash equilibrium. We will call these growth equilibria. Following the VNM utility theorem, we will show that for a class of capital games called positive capital games we can create a corresponding standard game where the payoffs are utilities, and the Nash equilibria of the created game correspond to the growth equilibria of the capital game. The conversion from positive capital games to standard games is shown to be reversible, so their equilibria coincide exactly. Thus, for games where players use strategies that maximize the time-average growth rate of their capital, we inherit powerful tools, and complexity, from the existing theory of games.
The fundamental goals of this paper are (1) to illustrate clearly for readers that desirable observables should not necessarily be treated as VNM utilities, and (2) demonstrate how growth rates of stochastic processes can be used to compute VNM utility functions and reason about equilibria in games. Ultimately, we argue that many applications of game theory to real world problems should be modeled first as capital games, and any mapping from payoff observables over time to timeless utility functions should be explicitly stated. The assumed dynamics behind the construction of VNM utility functions should always be made explicit, because they are not implied by the definition of standard games. This prevents problems in mechanism design and operations research, and particularly the design of auctions, from ending up FUC’d.
2 Games and Nash Equilibrium
We begin by providing a background on concepts, definitions, and key results in the Game Theory literature.
Definition 2 (Standard Game).
A standard game is a finite, n-person game characterized (in normal form) by a tuple $(N,A,u)$ where:
  • N is a finite set of $n\geq 1$ players, indexed by $i$;
  • $A=(A_{1},\ldots,A_{n})$ where $A_{i}$ is a finite set of actions (or pure strategies) available to player $i$
  • $u=(u_{1},\ldots,u_{n})$ where $u_{i}:A_{1}\times\ldots\times A_{n}\rightarrow\mathbb{R}$ is a real-valued function that determines the utility of every player based on the action profile
Action Profiles and Payoffs
We let $\mathcal{A}=A_{1}\times\ldots\times A_{n}$ denote the set of all possible action profiles, where each action profile $a=(a_{1},\ldots,a_{n})\in\mathcal{A}$ specifies a single action for every player. We can see by the definition of $u_{i}$ that the utility of each player depends on the collective actions of all players, not just their own (when $n>1$).
Mixed Strategy Profiles
Players may use mixed strategies, which are probability distributions over their available actions. For any set $X$, let $\Delta(X)$ denote the set of all probability distributions over $X$. Then the set of strategies $S_{i}$ available to player $i$ is $S_{i}=\Delta(A_{i})$. Players select their strategies simultaneously and independently.
A mixed strategy profile is given by $s=(s_{1},\ldots,s_{n})\in\mathcal{S}=S_{1}\times\ldots\times S_{n}$. We denote by $s_{i}(a^{j})$ the probability that strategy $s_{i}$ assigns to taking action $a^{j}\in A_{i}$. Given a mixed strategy profile $s$, we define $s(a)=\prod_{i\in N}s_{i}(a_{i})$ to be the probability of the action profile $a=(a_{1},\ldots,a_{n})$ being realized under $s$. The support of $s_{i}$ is all actions $a^{j}\in A_{i}$ such that $s_{i}(a^{j})>0$, and similarly, the support of $s$ is all action profiles $a$ such that $s(a)>0$.
All players are assumed to have complete and perfect information, so they know the strategy spaces and utility functions of the other players, and it is common knowledge that all players have complete and perfect information.
Definition 3 (Expected Utility).
Given normal-form game $(N,A,u)$, and a mixed strategy profile $s=(s_{1},\ldots,s_{n})$, the expected utility of player $i$ is defined as:
$E[u_{i}|s]=\sum\limits_{a\in\mathcal{A}}u_{i}(a)s(a)$
[…]
Definition 4 (Response).
A response is a choice of strategy $s_{i}$ given a profile $s_{-i}$ of the strategies of all other players.
Definition 5 (Best Response).
Player $i$’s best response to strategy profile $s_{-i}$ is a mixed strategy $\bar{s}_{i}\in S_{i}$ such that $E[u_{i}|(\bar{s}_{i},s_{-i})]\geq E[u_{i}|(s_{i},s_{-i})]$ for all $s_{i}\in S_{i}$.
A Nash equilibrium is a strategy profile such that no player $i$ can increase their expected utility by unilaterally changing their strategy. In other words, $s$ is a Nash equilibrium if for all players $i$, $s_{i}$ is a best response to $s_{-i}$.
Definition 6 (Nash Equilibrium).
A Nash equilibrium is a strategy profile $s$ such that $s_{i}$ is a best response to $s_{-i}$ for all players $i\in N$, maximizing their expected utility.
Theorem 2 (Existence of Nash Equilibria).
All standard games have at least one Nash equilibrium (Nash, 1950).
Although Nash equilibria always exist, computing an equilibrium of a standard game is known to be PPAD-complete (Daskalakis et al., 2009).
3 Gambles and Best Responses
We will introduce a core definition from Ergodicity Economics called a gamble, which we will apply to the study of games.
Definition 7 (Gamble).
A gamble is a tuple $(Q,\delta t)$ where $Q$ is a discrete random variable that takes one of $K$ real values $\{q_{1},\ldots,q_{K}\}$ each with probability $\{p_{1},\ldots,p_{K}\}$, and $\delta t$ is a unit of time called duration.
A gamble is a lottery whose outcomes are real-valued and which takes place over some amount of time.
Definition 8 (Gamble Problem).
A gamble problem is a problem of choosing between two or more available gambles.
Proposition 1.
In a standard game, the problem of a player choosing their best response $s_{i}\in S_{i}$ given $s_{-i}$ and $u_{i}$ is a gamble problem if we let $\delta t$ be the duration of the game.
Proof.
Given $s_{-i}$, each choice of strategy $s_{i}\in S_{i}$ creates a different strategy profile $s=(s_{i},s_{-i})$, which assigns a probability $s(a)$ to each possible action profile $a$, of which there are finitely many. Since $u_{i}$ is deterministic, this means that $s$ also induces a probability distribution over the finite number of possible utilities of player $i$. Of course the utilities of $i$ are real-valued, and $\sum_{a\in A}s_{i}(a)=1$. Therefore choosing a strategy $s_{i}$ given $s_{-i}$ means choosing a gamble defined by possible realizations $\{u_{i}(a)\}_{a\in\mathcal{A}}$ with associated probabilities $\{s(a)\}_{a\in\mathcal{A}}$, and $\delta t$ is the duration of the game. ∎
From the definition of a VNM utility function, we have built in the assumption that the player wishes to maximize the expected value of utility. In practice, we are given gambles with the values $\{q_{1},\ldots,q_{K}\}$, but we cannot assume these values are the same units as our utility.
As Peters and Adamou (2025) argue, the definition of a gamble on its own does not provide sufficient information to choose between gambles. This means the definition of a gamble does not given us enough information to determine our VNM utility function. We cannot generally compare possible gambles without (1) knowing how the outcome affects the future, e.g., how the outcome impacts future decisions; (2) our circumstances, i.e., in the relevant units ($q_{i}$), how much are we initially endowed with before the gamble; and (3) our personal preferences, e.g., risk tolerance. We will put aside the psychological third aspect for our purposes, and extend our treatment of games based on the first two; by having players start with some initial endowment of an observable of interest (i.e., capital) and treat each game as a single time step in a (potentially infinite) sequence of (potentially independent) decisions each player faces, from which we will derive our VNM utility function.
In practice, each outcome of a gamble could have a separate duration. For brevity we will assume all outcomes of a gamble have the same duration.
4 Capital Games
We now define capital games, which are like standard games except that (1) payoffs are in arbitrary units of “capital” rather than units of “utility”, (2) every player has some initial capital endowment, (3) every player’s payoffs are characterized by dynamics which take place over some amount of time.
Definition 9 (Capital Game).
A finite capital game is characterized by a tuple $(N,W,A,x,D,f)$ where
  • N is a finite set of $n\geq 1$ players, indexed by $i$;
  • $A=(A_{1},\ldots,A_{n})$ where $A_{i}$ is a finite set of actions available to player $i$
  • $x=(x_{1},\ldots,x_{n})$ where $x_{i}:A_{1}\times\ldots\times A_{n}\rightarrow\mathbb{R}$ is a real-valued function that determines the payoff of every player based on the action profile in units of capital
  • $W=(w_{1},\ldots,w_{n})\in\mathbb{R}^{n}$ is an initial endowment for each player, in units of capital.
  • $D=(\delta t_{1},\ldots,\delta t_{n})$ the duration of the game for each player
  • $f=(f_{1},\ldots,f_{n})$ where the $f_{i}$ are each deterministic functions that captures player $i$’s capital dynamics $f_{i}(x_{i}(a),w_{i},\delta t_{i})\in\mathbb{R}$.
Definition 10 (Dynamics Linearization).
Function $v_{i}$ is a linearization of capital dynamics $f_{i}$ if $v_{i}(f_{i}(x_{i}(a),w_{i},\delta t_{i}))=\frac{v_{i}(x_{i}(a))-v_{i}(w_{i})}{\delta t_{i}}$ for all $a\in\mathcal{A}$.
For brevity we will denote capital games by $(N,W,A,x,f)$ and assume the game duration is equal for all players who act simultaneously and receive their payoffs simultaneously, so the duration is a constant normalized to 1. We will therefore also use the shorthand $f_{i}(x_{i}(a),w_{i})$ for capital dynamics.
We will assume that the capital dynamics of all players are linearizable, meaning that a dynamics linearization exists for each of them. Many different linearizable capital dynamics are possible. The two simplest dynamics are additive dynamics $f_{i}(x_{i}(a),w_{i})=x_{i}(a)-w_{i}$ and multiplicative dynamics $f_{i}(x_{i}(a),w_{i})=\frac{x_{i}(a)}{w_{i}}$. For additive dynamics $v_{i}(x)=x$ is a linearization because $f_{i}$ is already a linear function. For multiplicative dynamics, $v_{i}(x)=\ln x$ is a linearization because $\ln\frac{x_{i}(a)}{w_{i}}=\ln x_{i}(a)-\ln w_{i}$, which is defined for all positive capital games.
Definition 11 (Positive Capital Game).
A capital game $(N,W,A,x,f)$ is positive if $w_{i}>0$ and $x_{i}(a)>0$ for players $i$ and action profiles $a$.
5 Best Responses and Growth Equilibria
Players determine their best responses in accordance with the decision axiom, using all available information.
Proposition 2 (Decision Axiom (Peters and Adamou, 2025)).
Players seek to maximize the time-average growth rate $\bar{g}_{i}$ of their capital.
The time-average growth rate of capital is ergodic, and in our setting
$\bar{g}_{i}=E[v_{i}(f_{i}(x_{i}(a),w_{i},\delta t_{i}))|s]$
Definition 12 (Best Response).
Given a capital game $(N,W,A,x,f)$ and opponents’ strategy profile $s_{-i}$, the best response of player $i$ is the strategy $s_{i}$ that maximizes the time-average growth rate of player $i$’s capital:
$\bar{s}_{i}=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}E[v_{i}(f_{i}(x_{i}(a),w_{i},\delta t_{i}))|s]$
The time-average growth rate of player $i$ in a game depends on the full strategy profile $s$. We can define a general growth equilibrium $s^{*}$ when all players are playing their best responses, maximizing the time-average growth of their capital conditioned on the strategies of the other players.
Definition 13 (Growth Equilibrium).
Given an capital game $(N,W,A,x,f)$, a growth equilibrium is a strategy profile $s^{*}=(s^{*}_{i},s^{*}_{-i})$ such that $s^{*}_{i}$ is a best response to $s^{*}_{-i}$ for all players $i\in N$.
Recall that we assume complete information, so each player must know the capital dynamics and duration of the other players to infer their equilibrium strategies.
Capital dynamics can be inhomogenous, or different for every player. Consider a gambler playing roulette in a casino. The gambler controls what fraction of their endowment they bet in each roll of the roulette wheel. Suppose the gambler bets a fixed fraction of their wealth on every game they play. Then the player’s wealth faces multiplicative dynamics, but the casino’s wealth experiences a different dynamics because the bet sizes are not a fixed fraction of the casino’s wealth.
We prove that for all positive capital games where each player’s capital dynamics admit a linearization, there is a correspondence between their growth equilibria and the Nash equilibria of a standard game, assuming players’ capital dynamics are common knowledge.
Theorem 3 (Equilibrium Correspondence).
Let $G=(N,W,A,x,f,D)$ be a positive capital game, where $\delta t_{i}=1$ and capital dynamics $(f_{1},\ldots,f_{n})$ can be linearized by $(v_{1},\ldots,v_{n})$, for all $i\in N,a\in\mathcal{A}$. Let $G^{\prime}=(N,A,u)$ be the standard game where $u_{i}(a)=v_{i}(f_{i}(x_{i}(a),w_{i}))$ for all $i\in N,a\in\mathcal{A}$. Then the Nash equilibria of $G^{\prime}$ are exactly the growth equilibria of $G$.
Proof.
Given $s_{-i}$, the set of best responses of player $i$ is
$\bar{S}_{i}$ $=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}v_{i}(f_{i}(x_{i}(a),w_{i}))s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}u_{i}(a)s(a)=\bar{S}^{\prime}_{i}$
In a growth equilibrium $s^{*}$ of $G$, all players are playing their best responses to one another under their respective capital dynamics, which means that in $G^{\prime}$ all players are playing their best responses to one another in $s^{*}$, and $s^{*}$ is therefore a Nash equilibrium. ∎
Corollary 1 (Existence of Growth Equilibria).
Every positive capital game with linearizable dynamics has at least one growth equilibrium.
Corollary 2 (Complexity of Computing Growth Equilibria).
Computing a growth equilibrium for positive capital games is PPAD-complete.
6 Additive and Multiplicative Dynamics
For any positive capital game, we can use dynamics linearization to create a standard game whose Nash equilibria are exactly the growth equilibria in our original game. In general, this does not require the players to have the same capital dynamics or the same linearization functions. But here, we will show the construction and correspondence in detail when all players have additive and multiplicative dynamics.
6.1 Additive Dynamics and Nash Equilibria
Let’s first look at the case where all players have additive capital dynamics. Under additive capital dynamics, $v_{i}(x)=x$ is a linearization because $f_{i}$ is already linear. That means capital is utility and utility is capital. Every player’s best response is to maximize the expected value of their capital. With additive dynamics, the growth equilibrium is the Nash equilibrium.
Given $s_{-i}$, each best response is
$\bar{s}_{i}$ $=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}v_{i}(f_{i}(x_{i}(a),w_{i}))s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}(v_{i}(x_{i}(a))-v_{i}(w_{i}))s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}(x_{i}(a)-w_{i})s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}x_{i}(a)s(a)$
Given a positive capital game $G$, we can construct a standard game $G^{\prime}=(N,A,u)$ by letting $u_{i}(a)=x_{i}(a)-w_{i}$ for all players and action profiles. Note that the players, action spaces, and strategy spaces are the same as in the capital game. We can see immediately that the best responses in the capital game are also the best responses in the standard game:
$\bar{s}^{\prime}_{i}=\bar{s}_{i}$ $=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}u_{i}(a)s(a)$
When all players respond this way, the growth equilibria in the capital game $G$ are the Nash equilibria in the standard game $G^{\prime}$.
This argument is reversible. Given a standard game $G^{\prime}$ we can construct a positive positive capital game $G$ by selecting $w_{i}>0$ for each player, $x_{i}(a)=u_{i}(a)+w_{i}$, and $f_{i}(x_{i}(a),w_{i})=x_{i}(a)-w_{i}$ for all players and action profiles. Formally, we must further specify $\delta t_{i}=1$ for all players. We can now follow the steps of the argument above in reverse to see that the growth equilibria of $G$ are the Nash equilibria of $G^{\prime}$.
Notice that for any standard game with additive dynamics we can set the endowments of each player to be whatever we want in the capital game. Thus, we can choose them to be the same as any positive capital game of interest. Therefore, there is a one-to-one correspondence between growth equilibria in positive capital games with additive dynamics for all players and Nash equilibria in the corresponding games.
6.2 Multiplicative Dynamics and Kelly Equilibria
Let’s look at the case where all players have multiplicative capital dynamics with linearization $v(x)=\ln x$. Player $i$ chooses their best response based on maximizing their time-average growth rate under multiplicative dynamics. Given $s_{-i}$, each best response is
$\bar{s}_{i}$ $=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}v_{i}(f_{i}(x_{i}(a),w_{i}))s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}(v_{i}(x_{i}(a))-v_{i}(w_{i}))s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}(\ln(x_{i}(a))-\ln(w_{i}))s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}\ln(x_{i}(a))s(a)$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\sum\limits_{a\in\mathcal{A}}\ln(x_{i}(a)^{s(a)})$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\ln\prod\limits_{a\in\mathcal{A}}x_{i}(a)^{s(a)}$
$=\operatorname*{arg\,max}\limits_{s_{i}\in S_{i}}\prod\limits_{a\in\mathcal{A}}x_{i}(a)^{s(a)}$
which is well-defined because $\ln(x)$ is defined for all $x>0$, and we are considering positive capital games.
Given a positive capital game $G=(N,W,A,x,f)$ where all players have multiplicative capital dynamics (and all $\delta t_{i}=1$), we can create a standard game $G^{\prime}=(N,A,u)$ where $u_{i}(a)=\ln x_{i}(a)-\ln w_{i}$ for all players $i$ and action profiles $a$. As with additive dynamics, the players, action spaces, and strategy spaces are the same, and now the growth equilibria of $G$ are all Nash equilibria in $G^{\prime}$.
Once again we can reveres the process. Given a standard game $G^{\prime}=(N,A,u)$, create a positive capital game $G=(N,W,A,x,f)$ where $w_{i}>0$, $x_{i}(a)=e^{u_{i}(a)}w_{i}$, $f_{i}(x_{i}(a),w_{i})=\frac{x_{i}(a)}{w_{i}}$, and $\delta t_{i}=1$, for all players and action profiles. The capital game is guaranteed to be positive because $e^{x}>0$ for all $x\in\mathbb{R}$. We can follow the steps above in reverse to see that the growth equilibria of $G$ are the Nash equilibria of $G^{\prime}$.
7 Pure Growth Equilibria
Just as we can define pure Nash equilibria, we can define pure growth equilibria where players are restricted to pure strategies rather than mixed strategies.
Definition 14 (Growth Rate).
Given a monotonically increasing function $v$, and action profile $a$, we define the growth rate of the utility of player $i$ to be $g^{v}_{i}(a)=\frac{v(x_{i}(a))-v(w_{i})}{\delta t_{i}}$.
Definition 15 (Best Response in Pure Strategies).
When all players are restricted to pure strategies (i.e., actions), a best response to $a_{-i}$ is
$a^{*}_{i}=\operatorname*{arg\,max}\limits_{a_{i}\in A_{i}}g^{v}_{i}(a_{i},a_{-i})$
which does not depend on the choice of function $v$.
Definition 16 (Pure Growth Equilibrium).
A pure growth equilibrium is a strategy profile, which is also an action profile, $a^{*}$ such that every player $i$ is playing a best response in pure strategies $a_{i}^{*}$ to $a^{*}_{-i}$.
Theorem 4.
If $a^{*}$ is a pure growth equilibrium for any capital dynamics, then it is also a pure growth equilibrium for all other capital dynamics.
Proof.
If there exists a pure growth equilibrium $a^{*}$, then each strategy $a_{i}$ is a choice between actions, rather than probability distributions, and each player $i$ is choosing an action deterministically that maximizes their payoff. This maximizes their deterministic growth rate, and therefore maximizes their time-average growth rate under any assumed dynamics. ∎
Preferences over degenerate, deterministic lotteries are exactly preferences over outcomes, so the observable (capital) is a VNM utility. The fallacy of utility conflation does not appear in deterministic settings, because any (strictly) monotonic function $v$ preserves the ordering of a set of real values.
8 Deriving Dynamics
In a non-cooperative capital game, a best response is one that maximizes time-average growth rate of capital. To translate from capital to VNM utility we need linearization of dynamics. But where to these dynamics come from?
Let’s look back at the infamous coin flip from Peters and Gell-Mann (2016). Consider a player with endowment $w=100$, action space $A=\{a_{1},a_{2}\}$, and payoffs $x(a^{1})=150$ and $x(a^{2})=60$. If dynamics are additive, then $f(x(a_{1}),w)=150-100=50$ and $f(x(a_{2}),w)=60-100=-40$, and the corresponding standard game has $u_{i}(a)=f(x(a),w)$, so the utilities are $50$ and 40-40. On the other hand, if the dynamics are multiplicative, then $f(x(a_{1}),w)=\frac{150}{100}=1.5$ and $f(x(a_{2}),w)=\frac{60}{100}=0.6$, and the corresponding standard game has $u_{i}(a)=\ln f(x(a),w)$, so the utilities are $\ln 1.5\approx 0.405$ and $\ln 0.6\approx-0.511$. The key observation here is that the payoffs and endowments together do not imply the dynamics, and the dynamics must be specified in order to derive the VNM utilities for players who wish to maximize the time-average growth rate of their capital.
So why should a player have additive or multiplicative dynamics or a totally different dynamics altogether? In standard games, the meaning of utilities could be ignored because regardless of what it represents the player seeks to maximize its expected value. This is built into the definition of a utility. But in capital games we cannot make this assumption. The dynamics depend on factors not otherwise captured in the definition of a capital game. The answer is application-specific.
Dynamics from Game Sequences
One setting in which dynamics can be derived is in sequences of games where a player’s payoff in each game becomes their endowment in the next, and their dynamics stay constant while the payoffs change in each game. For example, if the coin flip game is repeated many times in sequence with multiplicative dynamics, then the payoffs keep changing in every round but the ratio between payoff and endowment stays the same for every action profile. In general, dynamics may be uncertain, but this is not captured in our model with deterministic dynamics. With deterministic dynamics, the only source of uncertainty comes from players’ mixed strategies.